# Towards the Worst-case Robustness of Large Language Models

### v3

@@ 5 Case Studies
In this section, we conduct two case studies, analyzing the certified robustness using two popular smoothing kernel $p(\bm{z}|\bm{x})$ – a uniform kernel (i.e., the forward distribution in diffusion models [54, 20]) and an absorbing kernel (i.e., the forward distribution in mask generation [55, 56]). We show that when they achieve the same standard accuracy, the robustness of the former is strictly greater than that of the latter (and they are equal when the vocabulary size $|\mathcal{V}|\to\infty$).
@@ 5.1 Certified Robustness on Absorbing Kernel
(Absorbing Kernel). An absorbing kernel perturbs each token independently. Each token is replaced with a special masked token $[M]$ with probability $\beta$, and remains unchanged with probability $\bar{\beta}=1-\beta$:
$p(\bm{z}_{i}|\bm{x}_{i})=\begin{cases}\bm{x}_{i}&\text{w.p.\ }\beta,\\ \text{[M]}&\text{w.p.\ }\bar{\beta}=1-\beta.\end{cases}$ (2)
For simplicity, we use the subscript $p$ to denote the common part of $\bm{x}$ and $\bm{x}_{adv}$, and the subscript $s$ to denote the differing part of $\bm{x}$ and $\bm{x}_{adv}$. Clearly, $\bm{x}_{p}=\bm{x}_{adv_{p}}$.
@@
(c) Different vocabulary size.
@@
To apply fractional knapsack solvers to specific smoothing kernels, a brute-force approach is to enumerate all possible $\bm{z}$ and perform Algorithm 1 for each $\bm{z}$. However, fractional knapsack solvers only depend on the value-to-weight ratio and the total weight of items with a given value-to-weight ratio. If multiple items share the same value-to-weight ratio, we can group these items into categories and calculate the total weight (volume) for each category. Formally, the volume $v(\gamma)$ for a trading rate $\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}=\gamma$ is defined as:
@@
(Proof in Sec. D.2) Divide $\mathcal{V}^{N}$ into $L_{1}$ and $L_{2}$ that $L_{1}\cup L_{2}=\mathcal{V}^{N}$ and $L_{1}\cap L_{2}=\emptyset$, where $L_{1}=\{\bm{z}\in\mathcal{V}^{N}\mid\bm{z}_{s}\text{ are all masked tokens}\}$, $L_{2}=\{\bm{z}\in\mathcal{V}^{N}\mid\bm{z}_{s}\text{ are not all masked tokens}\}$. Clearly, we have the trading rate:
@@
By applying these results to Algorithm 1, we show that for the absorbing kernel, if $p_{A}=g(\bm{x})\leq 1-\beta^{d}$, no robustness guarantee can be obtained. For $p_{A}\geq 1-\beta^{d}$, we can obtain a robustness guarantee that $p_{adv}=g(\bm{x}_{adv})\geq p_{A}-(1-\beta^{d})$, with a maximum of $\beta^{d}$, as illustrated in Figure 1(a).
Table 2: The average certified $\ell_{0}$ radius against $\ell_{0}$ attack.
Vicuna-7B Llama2-7B Llama3-8B GPT-4o Human
Absorb 1.00 1.92 1.82 2.00 2.12
Uniform 1.02 1.86 1.54 2.02 2.12
SmoothLLM 2.25 3.24 3.16 3.84 4.04
@@ 5.2 Certified Robustness on Uniform Kernel
A notable property of the uniform kernel is that if $g(\bm{x})=1$, then $g(\bm{x}_{adv})$ is also one. This occurs because the support of $p(\bm{z}|\bm{x})$ spans the entire space $\mathcal{V}^{N}$. When $g(\bm{x})=\sum_{\bm{z}}f(\bm{z})p(\bm{z}|\bm{x})=1$, it implies that $f(\bm{z})=1$ for all $\bm{z}$. Consequently, $g(\bm{x}_{adv})=\sum_{\bm{z}}f(\bm{z})p(\bm{z}|\bm{x}_{adv})$ will also equal 1. In contrast, with the absorbing kernel, $g(\bm{x}_{adv})$ cannot exceed $\beta^{d}$. From this perspective, the uniform kernel closely resembles the behavior of the Gaussian distribution in the image domain, where the certified radius can also potentially be infinite [23, 45].
@@
(Proof in Sec. D.4.) The certified radius of the uniform kernel is always greater than or equal to that of the absorbing kernel given the same accuracy $p_{A}$, threshold $\tau$, and perturbation probability $\beta$, i.e.,
@@ 6 Experiment
In this section, we demonstrate the experimental details of both empirical evaluation for upper bounding worst-case robustness and theoretical certification for lower bounding worst-case robustness.
@@ 6.1 Empirical Evaluations
Settings. We conduct both black-box evaluations to demonstrate practical usage (Sec. G.3) and white-box evaluations (Table 1) to establish the upper bound of worst-case robustness. Following [10, 17, 57], we use the AdvBench dataset [10]. We perform suffix attacks that append $d=20$ adversarial tokens as a suffix to the original request and optimize these appended tokens. We set $\beta=0.25$. Refer to Appendix G for other details.
Results. As shown in Sec. G.3, all defenses achieve reasonable performance in black-box settings, demonstrating their high practicality. For white-box settings, see Sec. 3 for details.
@@ 6.2 Certified Robustness
Settings. We use the AdvBench dataset [10] to evaluate certified lower bounds for three previous empirical defenses: uniform kernel [30, 20], absorbing kernel [56, 55, 29], and SmoothLLM [31] (i.e., a uniform kernel applied to each character instead of each token). Note that the results for SmoothLLM presented in this paper certify character-level robustness rather than token-level robustness. We welcome contributions of new stochastic defenses for large language models and will provide their lower bounds using our solver, incorporating results into the main table.
We focus on certifying safety against two types of attacks. In the $\ell_{0}$ attack, we set $\beta=0.1$ and apply these defenses to the entire sentence, thereby certifying the $\ell_{0}$ radius. In the Suffix Attack, we set $\beta=0.25$, pad the input sentence with 50 arbitrary tokens, and apply these defenses to all tokens except the first $k$ tokens. Safety detectors are constructed by adjusting the prompt of the LLM (see Sec. G.1). This prompt is highly conservative, ensuring a 0% false positive rate on normal requests across datasets, including [58, 59, 60, 61].
Baseline. For $\ell_{0}$ attacks, certified radii cannot be arbitrarily large. For example, "how to make a bomb" can become "how to make a cake" by changing one token, thus the certified radius of this sentence cannot exceed 0. The “Human” baseline serves as an upper bound for the certified radius; see Sec. I.2 for details. For suffix attacks, we compare randomized smoothing with the method of Kumar et al. [62], which deletes the suffix and evaluates the detector on the resulting sentence. Consequently, the certified robustness equals the clean accuracy (i.e., 1), and the certified radius is infinite. All randomized smoothing methods degrade to Kumar et al. [62] when $\beta\to 1$.
Results. As shown in Table 2, for $\ell_{0}$ attacks, we achieve a certified radius of $2.02$. The better the base model, the higher the true positive rate, and thus, the higher the certified radius. For the AdvBench dataset, the obtained theoretical lower bound is close to the human performance. However, this does not always hold true, especially for datasets containing longer requests (Sec. G.7). This may require a fundamental improvement on the randomized smoothing paradigm, e.g., relying on more neural network-dependent variables rather than a single $p_{A}$. For adversarial suffix attacks, we achieve an average certified radius of $6.41$ (with $\beta=0.25$), while practical settings focus on suffix lengths of $20$. This demonstrates that it is relatively easy to obtain a certified radius with strong practical significance in the suffix attack settings due to its simplicity [62].
Table 3: The average certified length against suffix attack using Llama-3-8B [3].
@@
Kumar et al. [62] $\infty$ $\infty$ $\infty$ $\infty$
Smoothness-utility Trade-off. As $\beta$ approaches 1, the distribution of diffused samples becomes identical for both benign and adversarial inputs. In this scenario, the base model cannot distinguish whether noisy examples originate from benign or adversarial inputs. Consequently, the function $g$ becomes overly smooth, producing a constant output regardless of the input. Since we require a false positive rate of 0, in the $\ell_{0}$ attack setting, this directly results in a certified radius of 0. In the suffix attack setting, the detector relies solely on the prefix, leading to a certified radius of either 0 or $\infty$, and all randomized smoothing methods degrade to Kumar et al. [62]. A certified radius of $\infty$ may be undesirable, as adding a few tokens can significantly alter the semantics of an input sentence (see Sec. H.3). Typically, we choose $\beta=0.25$, as this value avoids masking critical information in the sentence and prevents oversmoothing.
@@ 7 Conclusion
In this work, we investigate the worst-case robustness of large language models. We upper bound the worst-case robustness of previous defenses by proposing a strong adaptive attack that strictly ensures the consistency in tokenization between optimization and inference. We also lower bound the worst-case robustness of all randomization-based defenses by reducing the functional optimization to a fractal knapsack problem or 0-1 knapsack problem. We conduct two case studies on smoothing the distribution of the diffusion models and masked generation, analyze their certified lower bound and clean accuracy, demonstrating their relationship. We also provide theoretical analysis on the relationship between certified robustness, smoothing distribution, and vocabulary size, and upper bound the certified lower bound by Bayesian error, offering insights into the upper limits of certified methods. See Appendix K for key takeaways.
@@ Appendix A Notations
$-\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}$ Value-to-weight ratio.
@@
$N$ (maximum) Input length.
@@ B.1 More Related Work on Jailbreak Attacks and Defenses
The jailbreak attack on LLMs primarily refers to inducing LLMs into generating harmful content that is unsafe or toxic to society [63, 64]. To achieve this goal, malicious attackers can craft jailbreaking prompts through manual design, optimization, or train a generative model. Manual-designed jailbreak prompts leverage heuristic perspectives like data distribution [36, 65, 12], psychology insights [66, 67, 68, 69] or cipher encoding [70, 71] to achieve this goal. Optimization-based attacks extend from manually designing by optimizing an adversarial prompt with certain loss functions, where they can optimize a prefix or suffix [10, 39, 17, 72, 73], or directly refine the jailbreaking prompt [74, 75, 76, 11, 77]. Besides, a thread of work toward fitting the jailbreak prompt distribution with a generative model [57, 78, 79, 80], effectively increasing the attack efficiency. Notably, there are also fine-tuning-based attacks that directly manipulate the alignment instead of designing prompts [81, 82, 83], posing another safety threat to LLMs.
From the defense perspective, various methods are proposed at different stages of generation. Pre-processing defenses are designed to detect potential jailbreaking prompts, typically aimed at adversarial suffix-based attacks that cause significantly high perplexity [48, 84]. Besides, prompt-based defenses add safety tokens during generation, which are manually designed [36, 85] or optimized [41, 86]. Finally, post-processing defenses detect jailbreaking with hidden spaces [87, 88] or toxicity detection [89, 90, 91].
@@ B.2 Adversarial Attacks and Defenses on Text Domain
Textual adversarial attacks [92, 93, 94] extend adversarial examples from vision space to discrete text space. Thus, a major challenge of textual attacks is the optimization process on discrete tokens, which include character, word, or sentence-level attacks. For instance, word-level attacks replace critical tokens with semantically similar alternatives to evade detection [55, 95], while character-level attacks insert misspellings or Unicode artifacts to bypass filters [96, 97]. Recent advances also employ generative models to automate the creation of adversarial examples [98, 99], producing fluent but malicious inputs that align with natural language patterns. These attacks highlight the vulnerability of text-based systems to carefully crafted inputs, even when perturbations are imperceptible to humans.
Defending against textual adversarial attacks also requires addressing the discrete nature of language. Adversarial training [100], which incorporates perturbed examples during model optimization, remains a cornerstone for improving the robustness of language models [101, 102]. A series of certified defenses with randomized smoothing techniques provide probabilistic guarantees against textual bounded perturbations [30, 103] was also proposed. The evolving landscape of text-domain adversarial robustness underscores the need for defenses that generalize across attack vectors while preserving linguistic integrity. However, these defenses and certifications are limited to conventional language models like sentence classifiers, yet the certified robustness of large generative models remains unexplored.

### v4

@@ 5 Case Studies
In this section, we conduct two case studies, analyzing the certified robustness on text data using two popular smoothing kernel $p(\bm{z}|\bm{x})$ – a uniform kernel (i.e., the forward distribution in diffusion models Meng et al. (2022); Lou et al. (2023)) and an absorbing kernel (i.e., the forward distribution in mask generation Jin et al. (2020); He et al. (2022)). We show that when they achieve the same standard accuracy, the robustness of the former is strictly greater than that of the latter (and they are equal when the vocabulary size $|\mathcal{V}|\to\infty$).
@@ 5.1 Certified Robustness on Absorbing Kernel
(Absorbing Kernel). We use the subscript $i$ to denote the $i$-th token of an input. An absorbing kernel perturbs each token $\bm{x}_{i}$ independently. Each token is replaced with a special masked token $\text{M}]$ with probability $\beta$, and remains unchanged with probability $\bar{\beta}=1-\beta$:
$p(\bm{z}_{i}|\bm{x}_{i})=\begin{cases}\bm{x}_{i}&\text{w.p.\ }\bar{\beta}=1-\beta,\\ [\text{M}]&\text{w.p.\ }\beta.\end{cases}$ (3)
For simplicity, let $P=\{i\mid\bm{x}_{i}=\bm{x}_{adv,i}\}$ denote the indices of common part between $\bm{x}$ and $\bm{x}_{adv}$, $S=\{i\mid\bm{x}_{i}\neq\bm{x}_{adv,i}\}$ denote the indices of differing part between $\bm{x}$ and $\bm{x}_{adv}$. We use subscripts $P$ and $S$ to denote the sets of tokens from the corresponding inputs, i.e., $\bm{x}_{P}=\{\bm{x}_{i}\mid i\in P\}$ and $\bm{x}_{S}=\{\bm{x}_{i}\mid i\in S\}$ [^3] This is a generalization of prefix/suffix in the context of LLM attacks..
@@
(c) Different $|\mathcal{V}|$ ($|\mathcal{V}_{1}|>|\mathcal{V}_{2}|$).
@@
To apply fractional knapsack solvers to specific smoothing kernels, a brute-force approach is to enumerate all possible $\bm{z}$ and perform Algorithm˜1 for each $\bm{z}$. However, fractional knapsack solvers only depend on the value-to-weight ratio and the total weight of items with a given value-to-weight ratio. If multiple items share the same value-to-weight ratio, we can group these items into categories and calculate the total weight (volume) for each category. Formally, the volume $v(\gamma)$ for a trading rate $\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}=\gamma$ is defined as:
@@
(Proof in Sec.˜D.2) Divide $\mathcal{V}^{N}$ into $L_{1}$ and $L_{2}$ that $L_{1}\cup L_{2}=\mathcal{V}^{N}$ and $L_{1}\cap L_{2}=\emptyset$, where $L_{1}=\{\bm{z}\in\mathcal{V}^{N}\mid\bm{z}_{S}\text{ are all masked tokens}\}$, $L_{2}=\{\bm{z}\in\mathcal{V}^{N}\mid\bm{z}_{S}\text{ are not all masked tokens}\}$. Clearly, we have the trading rate:
@@
By applying these results to Algorithm˜1, we show that for the absorbing kernel, if $p_{A}=g(\bm{x})\leq 1-\beta^{d}$, no robustness guarantee can be obtained. For $p_{A}\geq 1-\beta^{d}$, we can obtain a robustness guarantee that $p_{adv}=g(\bm{x}_{adv})\geq p_{A}-(1-\beta^{d})$, with a maximum of $\beta^{d}$, as illustrated in Figure˜1(a).
@@ 5.2 Certified Robustness on Uniform Kernel
A notable property of the uniform kernel is that if $g(\bm{x})=1$, then $g(\bm{x}_{adv})$ is also one. This occurs because the support of $p(\bm{z}|\bm{x})$ spans the entire space $\mathcal{V}^{N}$. When $g(\bm{x})=\sum_{\bm{z}}f(\bm{z})p(\bm{z}|\bm{x})=1$, it implies that $f(\bm{z})=1$ for all $\bm{z}$. Consequently, $g(\bm{x}_{adv})=\sum_{\bm{z}}f(\bm{z})p(\bm{z}|\bm{x}_{adv})$ will also equal 1. In contrast, with the absorbing kernel, $g(\bm{x}_{adv})$ cannot exceed $\beta^{d}$. From this perspective, the uniform kernel closely resembles the behavior of the Gaussian distribution in the image domain, where the certified radius can also potentially be infinite Cohen et al. (2019); Salman et al. (2019).
@@
(Proof in Sec.˜D.4.) The certified radius of the uniform kernel is always greater than or equal to that of the absorbing kernel given the same accuracy $p_{A}$, threshold $\tau$, and perturbation probability $\beta$, i.e.,
@@ 6.1 Empirical Evaluations
Settings. We conduct both black-box evaluations to demonstrate practical usage (Sec.˜G.3) and white-box evaluations (Table˜1) to establish the upper bound of worst-case robustness. Following Zou et al. (2023); Jia et al. (2024); Liao & Sun (2024), we use the AdvBench dataset Zou et al. (2023). We perform suffix attacks that append $d=20$ adversarial tokens as a suffix to the original request and optimize these appended tokens. We set $\beta=0.25$. Refer to Appendix˜G for other details.
Results. As shown in Sec.˜G.3, all defenses achieve reasonable performance in black-box settings, demonstrating their high practicality. For white-box settings, see Sec.˜3 for details.
@@ 6.2 Certified Robustness
Settings. We use the AdvBench dataset Zou et al. (2023) to evaluate certified lower bounds for three previous empirical defenses: uniform kernel Lou et al. (2023), absorbing kernel He et al. (2022); Jin et al. (2020); Zeng et al. (2023), and SmoothLLM Robey et al. (2023) (i.e., a uniform kernel applied to each character instead of each token). Note that the results for SmoothLLM presented in this paper certify character-level robustness rather than token-level robustness.
We focus on certifying safety against two types of attacks. In the $\ell_{0}$ attack, we set $\beta=0.1$ and apply these defenses to the entire sentence, thereby certifying the $\ell_{0}$ radius. In the Suffix Attack, we set $\beta=0.25$, pad the input sentence with 50 arbitrary tokens, and apply these defenses to all tokens except the first $k$ tokens. Safety detectors are constructed by adjusting the prompt of the LLM (see Sec.˜G.1). This prompt is highly conservative, ensuring a 0% FPR on normal requests across datasets Zheng et al. (2024a); Cobbe et al. (2021); Hendrycks et al. (2020); Lin et al. (2021).
Baseline. For $\ell_{0}$ attacks, certified radii cannot be arbitrarily large. For example, "how to make a bomb" can become "how to make a cake" by changing one token, thus the certified radius of this sentence cannot exceed 0. The “Human” baseline serves as an upper bound for the certified radius; see Sec.˜I.2 for details. For suffix attacks, we compare randomized smoothing with the method of Kumar et al. (2023), which deletes the suffix and evaluates the detector on the resulting sentence. Consequently, the certified robustness equals the clean accuracy (i.e., 1), and the certified radius is infinite. All randomized smoothing methods degrade to Kumar et al. (2023) when $\beta\to 1$.
Results. As shown in Table˜3, for $\ell_{0}$ attacks, we achieve a certified radius of $2.02$. The better the base model, the higher the true positive rate, and thus, the higher the certified radius. For the AdvBench dataset, the obtained theoretical lower bound is close to the human performance. However, this does not always hold true, especially for datasets containing longer requests (Sec.˜G.7). This may require a fundamental improvement on the randomized smoothing paradigm, e.g., relying on more neural network-dependent variables rather than a single $p_{A}$. For adversarial suffix attacks, we achieve an average certified radius of $6.41$ (with $\beta=0.25$), while practical settings focus on suffix lengths of $20$. This demonstrates that it is relatively easy to obtain a certified radius with strong practical significance in the suffix attack settings due to its simplicity Kumar et al. (2023).
Table 2: The average certified $\ell_{0}$ radius.
Absorb Uniform SmoothLLM
Vicuna-7B 1.00 1.02 2.25
Llama2-7B 1.92 1.86 3.24
Llama3-8B 1.82 1.54 3.16
GPT-4o 2.00 2.02 3.84
Human 2.12 2.12 4.04
Table 3: The average certified length against suffix attack using Llama-3-8B.
@@
Kumar et al. (2023) $\infty$ $\infty$ $\infty$ $\infty$
Smoothness-utility Trade-off. As $\beta$ approaches 1, the distribution of diffused samples becomes identical for both benign and adversarial inputs. In this case, the base model cannot distinguish whether noisy examples originate from benign or adversarial inputs. Consequently, $g$ becomes overly smooth, producing a constant output regardless of the input. Since we require a false positive rate of 0, in the $\ell_{0}$ setting, this directly results in a certified radius of 0. In the suffix setting, the detector relies solely on the prefix, leading to a certified radius of either 0 or $\infty$, and all smoothing kernels degrade to Kumar et al. (2023). A certified radius of $\infty$ may be undesirable, as adding a few tokens can significantly alter the semantics of inputs (see Sec.˜H.4). Typically, we choose $\beta=0.25$, as this value avoids masking critical information and prevents oversmoothing.
@@ 7 Conclusion
In this work, we investigate the worst-case robustness of large language models. We upper bound the worst-case robustness of previous defenses by proposing a strong adaptive attack that strictly ensures the consistency in tokenization between optimization and inference. We also lower bound the worst-case robustness of all randomization-based defenses by reducing the functional optimization to a fractal knapsack problem or 0-1 knapsack problem. We conduct two case studies on smoothing the distribution of the diffusion models and masked generation, analyze their certified lower bound and clean accuracy, demonstrating their relationship. We also provide theoretical analysis on the relationship between certified robustness, smoothing distribution, and vocabulary size, and upper bound the certified lower bound by Bayesian error, offering insights into the upper limits of certified methods. See Appendix˜K for key takeaways.
@@ Appendix A Notations
$\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}$ Value-to-weight ratio.
@@
NN (maximum) Input length.
@@ B.1 More Related Work on Jailbreak Attacks and Defenses
The jailbreak attack on LLMs primarily refers to inducing LLMs into generating harmful content that is unsafe or toxic to society Chao et al. (2024); Zhou et al. (2024b). To achieve this goal, malicious attackers can craft jailbreaking prompts through manual design, optimization, or train a generative model. Manual-designed jailbreak prompts leverage heuristic perspectives like data distribution Wei et al. (2023b); Deng et al. (2023); Wei et al. (2023a), psychology insights Shen et al. (2024b); Zeng et al. (2024); Shen et al. (2024a); Li et al. (2023b) or cipher encoding Yuan et al. (2023); Handa et al. (2024) to achieve this goal. Optimization-based attacks extend from manually designing by optimizing an adversarial prompt with certain loss functions, where they can optimize a prefix or suffix Zou et al. (2023); Liu et al. (2023); Jia et al. (2024); Zhang & Wei (2025); Li et al. (2024a), or directly refine the jailbreaking prompt Dong et al. (2023); Chen et al. (2024c); Zheng et al. (2024b); Chao et al. (2023); Liu et al. (2024a). Besides, a thread of work toward fitting the jailbreak prompt distribution with a generative model Liao & Sun (2024); Kumar et al. (2024); Paulus et al. (2024); Basani & Zhang (2024), effectively increasing the attack efficiency. Notably, there are also fine-tuning-based attacks that directly manipulate the alignment instead of designing prompts Qi et al. (2023); Yang et al. (2023); Zhang et al. (2024b), posing another safety threat to LLMs.
From the defense perspective, various methods are proposed at different stages of generation. Pre-processing defenses are designed to detect potential jailbreaking prompts, typically aimed at adversarial suffix-based attacks that cause significantly high perplexity Jain et al. (2023); Alon & Kamfonas (2024). Besides, prompt-based defenses add safety tokens during generation, which are manually designed Wei et al. (2023b); Xie et al. (2023) or optimized Mo et al. (2024); Zhou et al. (2024a). Finally, post-processing defenses detect jailbreaking with hidden spaces Li et al. (2025); Galinkin & Sablotny (2024) or toxicity detection Wang et al. (2023b); Hu et al. (2024); Wang et al. (2024).
@@ B.2 Adversarial Attacks and Defenses on Text Domain
Textual adversarial attacks Morris et al. (2020); Wang et al. (2019b); Han et al. (2022) extend adversarial examples from vision space to discrete text space. Thus, a major challenge of textual attacks is the optimization process on discrete tokens, which include character, word, or sentence-level attacks. For instance, word-level attacks replace critical tokens with semantically similar alternatives to evade detection Jin et al. (2020); Zang et al. (2019), while character-level attacks insert misspellings or Unicode artifacts to bypass filters Ebrahimi et al. (2018); Rocamora et al. (2024). Recent advances also employ generative models to automate the creation of adversarial examples Ren et al. (2020); Li et al. (2023a), producing fluent but malicious inputs that align with natural language patterns. These attacks highlight the vulnerability of text-based systems to carefully crafted inputs, even when perturbations are imperceptible to humans.
Defending against textual adversarial attacks also requires addressing the discrete nature of language. Adversarial training Xiao et al. (2018), which incorporates perturbed examples during model optimization, remains a cornerstone for improving the robustness of language models Wang et al. (2019a); Gao et al. (2023). A series of certified defenses with randomized smoothing techniques provide probabilistic guarantees against textual bounded perturbations Jia et al. (2019); Wang et al. (2021) was also proposed. The evolving landscape of text-domain adversarial robustness underscores the need for defenses that generalize across attack vectors while preserving linguistic integrity. However, these defenses and certifications are limited to conventional language models like sentence classifiers, yet the certified robustness of large generative models remains unexplored.